Selling Bonds. Lunar Vacations needs to raise $6,200,000 for its new project (a golf course on the moon). Astro Investment Bank will sell the bond for a commission of 2.8%. The market yield is currently 7.7% on a 20 year semi-annual bonds. If Lunar wants to issue a 6.4% semi-annual coupon bond, how many bonds will it need to sell to raise the $6,200,000? Assume that all bonds are issued at par value of $1,000
Answer #1
                  K = Nx2
Bond Price =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =20x2
Bond Price =∑ [(6.4*1000/200)/(1 + 7.7/200)^k]     +   1000/(1 + 7.7/200)^20x2
                   k=1
Bond Price = 868.42

Proceeds per bond = price*(1-commission%)

=868.42*(1-0.028)=844.104

number of bonds = amount to raise/Proceeds per bond = 6200000/844.104=7345.066